OpenStudy (anonymous):

What is the standard form of (1+2sqrt-3)/(1+sqrt-3)?

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }+\frac{ \sqrt{3} }{ 4 }i \]

OpenStudy (anonymous):

im not that confident in that answer but tell me if its right or not